Find the LCM and HCF of the following integers by applying the prime factorisation method.
 12, 15 and 21 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

12, 15 and 21
Writing the product of prime factors for all the three numbers, we get,
12=2×2×3
15=5×3
21=7×3
Therefore,
HCF(12,15,21) = 3
LCM(12,15,21) = 2× 2 × 3 × 5 ×7 = 420

​Thanks